# Exponential Family Variational Flow Matching for Tabular Data Generation

### v1

@@ 1 Introduction
In this paper, we consider the application of flow matching to the modeling of tabular data, a ubiquitous data modality in a range of domains including finance, healthcare, and marketing. Tabular data pose unique modeling challenges due to heterogeneous features, potential missing values, and varying scales (Borisov et al., 2024; Wang et al., 2024). While some diffusion-based approaches have been developed (Kotelnikov et al., 2023; Zhang et al., 2024; Shi et al., 2024), models for tabular data are less widespread than their counterparts for images and text. In this context, recent work on variational flow matching (Eijkelboom et al., 2024; Zaghen et al., 2025) presents a promising avenue for generating mixed continuous and discrete features. VFM frames flow matching as an inference problem, in which the goal is to learn a variational posterior over data points that can be associated with the current interpolation point. This presents an opportunity to model heterogeneous data, which can be represented in a conceptually straightforward manner by matching the variational distribution to the data type for each variable.
@@ 2.3 Flow Matching as a Variational Inference Problem
$=-\mathbb{E}_{t,x_{1},x}\bigl{[}\log q_{t}^{\theta}(x_{1}\mid x)\bigr{]}+\text{const.}$. (8)
@@ 3.2 Exponential Families for VFM
$-\mathbb{E}_{t,x_{1},x}\left[\left(\mu_{t}(x)-\mu_{t}^{\theta}(x)\right)\cdot\nabla_{\theta}\eta_{t}^{\theta}(x)\right],$ (15)
@@ 4.1 Experimental setup
CTGAN $77.74\pm 0.15$ $62.08\pm 0.08$ $76.97\pm 0.39$ $86.90\pm 0.22$ $96.27\pm 0.14$ $96.96\pm 0.17$ $82.82$ $5$
TVAE $98.17\pm 0.17$ $85.57\pm 0.34$ $58.19\pm 0.26$ $86.19\pm 0.48$ $97.20\pm 0.10$ $86.41\pm 0.17$ $85.29$ $4$
GOGGLE $50.68$ $68.89$ $86.95$ $90.88$ $88.81$ $86.41$ $78.77$ $7$
TabDDPM $96.36\pm 0.20$ $97.59\pm 0.36$ $88.55\pm 0.68$ $88.59\pm 0.17$ $97.93\pm 0.30$- $79.83$ $6$
TabSyn $\mathbf{99.52\pm 0.10}$ $99.26\pm 0.27$ $\mathbf{99.16\pm 0.22}$ $99.38\pm 0.27$ $98.47\pm 0.10$ $96.80\pm 0.25$ $98.67$ $2$
TabDiff $99.02\pm 0.20$ $98.49\pm 0.28$ $99.11\pm 0.34$ $\mathbf{99.47\pm 0.21}$ $98.06\pm 0.24$ $\mathbf{97.36\pm 0.17}$ $98.59$ $3$
TabbyFlow $99.43\pm 0.18$ $\mathbf{99.31\pm 0.24}$ $98.96\pm 0.10$ $98.27\pm 0.39$ $\mathbf{98.93\pm 0.15}$ $97.22\pm 0.12$ $\mathbf{98.69}$ $\mathbf{1}$
@@
CTGAN $30.80\pm 0.20$ $18.22\pm 0.17$ $31.80\pm 0.35$ $11.75\pm 0.20$ $34.80\pm 0.10$ $24.97\pm 0.29$ $25.39$ $6$
TVAE $38.87\pm 0.31$ $23.13\pm 0.11$ $19.78\pm 0.10$ $32.44\pm 0.35$ $28.45\pm 0.08$ $29.66\pm 0.21$ $28.72$ $5$
GOGGLE $8.80$ $14.38$ $9.79$ $9.88$ $19.87$ $2.03$ $10.79$ $7$
TabDDPM $47.05\pm 0.25$ $47.83\pm 0.35$ $47.79\pm 0.25$ $48.46\pm 0.42$ $\mathbf{56.92\pm 0.13}$- $41.34$ $4$
TabSyn $47.56\pm 0.22$ $48.00\pm 0.35$ $48.95\pm 0.28$ $\mathbf{48.03\pm 0.23}$ $55.84\pm 0.19$ $45.04\pm 0.34$ $48.90$ $3$
@@ 4.2 Data Fidelity and Downstream Performance
Real .$927\pm.000$$770\pm.005$$926\pm.001$$946\pm.001$$423\pm.003$$842\pm.002$806.806.806.806
CTGAN .$886\pm.002$$696\pm.005$$875\pm.009$$855\pm.006$$\mathbf{.902\pm.019}$$\mathbf{.880\pm.016}$849.849.849.849
TVAE .$878\pm.004$$724\pm.005$$871\pm.006$$887\pm.004$$770\pm.011$$861\pm.016$831.831.831.831
GOGGLE .$778\pm.012$$584\pm.005$$658\pm.052$$654\pm.024$$709\pm.025$$877\pm.002$710.710.710.710
TabDDPM .$907\pm.001$$758\pm.004$$918\pm.005$$935\pm.004$$592\pm.011$$486\pm 3.04$766.766.766.766
TabSyn .$909\pm.001$$\mathbf{.763\pm.002}$$914\pm.004$$\mathbf{.937\pm.002}$$763\pm.002$$862\pm.004$$\mathbf{.858}$
TabDiff .$\mathbf{.912\pm.002}$$763\pm.005$$\mathbf{.921\pm.004}$$936\pm.003$$5$013 .$866\pm.021$826.826.826.826
TabbyFlow .$902\pm.002$$761\pm.003$$910\pm.006$$932\pm.003$$746\pm.008$$870\pm.005$854.854.854.854
@@
We evaluate the practical utility of synthetic data through MLE, following established protocols Kotelnikov et al. (2023); Zhang et al. (2024); Shi et al. (2024). This metric assesses how well models trained on synthetic data perform on real data, using XGBoost (Chen & Guestrin, 2016) for both classification (measured by AUC) and regression tasks (measured by RMSE). The results in Table 8 show that while TabbyFlow performs slightly below TabSyn on MLE scores, it maintains competitive performance in all data sets. This presents an interesting contrast to earlier metrics, where TabbyFlow often outperformed TabSyn. The small performance gap suggests that MLE alone may not fully capture synthetic data quality, particularly given that TabbyFlow achieves comparable results with a substantially simpler architecture than TabSyn’s VAE-diffusion framework. On the other hand, we evaluate the performance of the models in preserving privacy as measured by DCR. The results, see in Table 7, show that Tabbyflow outperforms all other models on average, albeit marginally.
@@ Acknowledgments
This project was support by the Bosch Center for Artificial Intelligence. JWvdM additionally acknowledges support from the European Union Horizon Framework Programme (Grant agreement ID: 101120237).
@@ A.2 Exponential Family VFM as Bregman Divergence
$\mathcal{L}_{\text{VFM}}(\theta)$ $=-\mathbb{E}_{t,x_{1},x}\left[\eta_{t}^{\theta}(x)\cdot\tau(x_{1})-A(\eta^{\theta}_{t}(x))\right]+\text{const.}$. (37)
$=-\mathbb{E}_{t,x_{1},x}\left[\mu^{\theta}_{t}(x)\cdot\eta^{\theta}_{t}(x)-A(\eta^{\theta}_{t}(x))+(\tau(x_{1})-\mu^{\theta}_{t}(x))\cdot\eta^{\theta}_{t}(x)\right]+\text{const.}$. (38)
$=-\mathbb{E}_{t,x_{1},x}\left[A^{*}(\mu_{t}^{\theta}(x))+(\tau(x_{1})-\mu_{t}^{\theta}(x))\cdot\eta_{t}^{\theta}(x)\right]+\text{const.}$. (39)
$=-\mathbb{E}_{t,x_{1},x}\left[-D_{A^{*}}\left(\tau(x_{1}),\mu_{t}^{\theta}(x)\right)+A^{*}(\tau(x_{1}))\right]+\text{const.}$. (40)
$=\mathbb{E}_{t,x_{1},x}\left[D_{A^{*}}\left(\tau(x_{1}),\mu_{t}^{\theta}(x)\right)\right]+\text{const.}$. (41)
$=\mathbb{E}_{t,x}\left[D_{A^{*}}\left(\mathbb{E}_{p_{t}(x_{1}\mid x)}[\tau(x_{1})],\mu_{t}^{\theta}(x)\right)\right]+\text{const.}$. (42)
$=\mathbb{E}_{t,x}\left[D_{A^{*}}\left(\mu_{t}(x),\mu_{t}^{\theta}(x)\right)\right]+\text{const.}$. (43)

### v2

@@ 1 Introduction
In this paper, we consider the application of flow matching to the modeling of tabular data, a ubiquitous data modality in a range of domains including finance, healthcare, and marketing. Tabular data pose unique modeling challenges due to heterogeneous features, potential missing values, and varying scales (Borisov et al., 2024; Wang et al., 2024). While some diffusion-based approaches have been developed (Kotelnikov et al., 2023; Zhang et al., 2024; Shi et al., 2024), models for tabular data are less widespread than their counterparts for images and text. In this context, recent work on variational flow matching (Eijkelboom et al., 2024; Zaghen et al., 2025; Eijkelboom et al., 2025) presents a promising avenue for generating mixed continuous and discrete features. VFM frames flow matching as an inference problem, in which the goal is to learn a variational posterior over data points that can be associated with the current interpolation point. This presents an opportunity to model heterogeneous data, which can be represented in a conceptually straightforward manner by matching the variational distribution to the data type for each variable.
@@ 2.3 Flow Matching as a Variational Inference Problem
$=-\mathbb{E}_{t,x_{1},x}\bigl{[}\log q_{t}^{\theta}(x_{1}\mid x)\bigr{]}+\text{const.}$ (8)
@@ 3.2 Exponential Families for VFM
$\mathbb{E}_{t,x_{1},x}\left[\left(\mu_{t}(x)-\mu_{t}^{\theta}(x)\right)\cdot\nabla_{\theta}\eta_{t}^{\theta}(x)\right],$ (15)
@@ 4.1 Experimental setup
CTGAN $77.74\pm 0.15$ $62.08\pm 0.08$ $76.97\pm 0.39$ $86.90\pm 0.22$ $27\pm 0.14 96.96$$96\pm 0.17$ $82.82$ 55
TVAE $98.17\pm 0.17$ $85.57\pm 0.34$ $58.19\pm 0.26$ $86.19\pm 0.48$ $97.20\pm 0.10$ $86.41\pm 0.17$ $85.29$ 44
GOGGLE $50.68$ $68.89$ $86.95$ $90.88$ $88.81$ $86.41$ $78.77$ 77
TabDDPM $96.36\pm 0.20$ $97.59\pm 0.36$ $8.55\pm 0.68 88$$8.59\pm 0.17$ $97.93\pm 0.30$ - $79.83$ 66
[…]
@@
CTGAN $30.80\pm 0.20$ $18.22\pm 0.17$ $31.80\pm 0.35$ $11.75\pm 0.20$ $34.80\pm 0.10$ $24.97\pm 0.29$ $25.39$ 66
TVAE $38.87\pm 0.31$ $23.13\pm 0.11$ $19.78\pm 0.10$ $32.44\pm 0.35$ $28.45\pm 0.08$ $29.66\pm 0.21$ $28.72$ 55
GOGGLE $8.80$ $14.38$ $9.79$ $9.88$ $19.87$ $2.03$ $10.79$ 77
TabDDPM $47.05\pm 0.25$ $47.83\pm 0.35$ $47.79\pm 0.25$ $48.46\pm 0.42$ $\mathbf{56.92\pm 0.13}$ - $41.34$ 44
TabSyn $47.56\pm 0.22$ $48.00\pm 0.35$ $48.95\pm 0.28$ $\mathbf{48.03\pm 0.23}$ $55.84\pm 0.19$ $45.04\pm 0.34$ $48.90$ 33
@@ 4.2 Data Fidelity and Downstream Performance
Real $927\pm.000 .$$770\pm.005 .$$926\pm.001 .$$946\pm.001 .$$423\pm.003 .$$842\pm.002 .$806.806
CTGAN $886\pm.002 .$$696\pm.005 .$$875\pm.009 .$$855\pm.006 .$$\mathbf{.902\pm.019}$ $\mathbf{.880\pm.016}$ .849.849
TVAE $878\pm.004 .$$724\pm.005 .$$71\pm.006 .88$$87\pm.004$ $770\pm.011 .$$861\pm.016 .$831.831
GOGGLE $778\pm.012 .$$584\pm.005 .$$658\pm.052 .$$654\pm.024 .$$709\pm.025 .$$877\pm.002 .$710.710
TabDDPM $907\pm.001 .$$758\pm.004 .$$918\pm.005 .$$935\pm.004 .$$592\pm.011 .$$486\pm 3.04 .$766.766
TabSyn $909\pm.001 .$$\mathbf{.763\pm.002}$ $914\pm.004 .$$\mathbf{.937\pm.002}$ $763\pm.002 .$$862\pm.004 .$$\mathbf{.858}$
TabDiff $\mathbf{.912\pm.002}$ $763\pm.005 .$$\mathbf{.921\pm.004}$ $936\pm.003 .$$555\pm.013 .$$866\pm.021 .$826.826
TabbyFlow $902\pm.002 .$$761\pm.003 .$$910\pm.006 .$$932\pm.003 .$$746\pm.008 .$$870\pm.005 .$854.854
@@
We evaluate the practical utility of synthetic data through MLE, following established protocols (Kotelnikov et al., 2023; Zhang et al., 2024; Shi et al., 2024). This metric assesses how well models trained on synthetic data perform on real data, using XGBoost (Chen & Guestrin, 2016) for both classification (measured by AUC) and regression tasks (measured by RMSE). The results in Table 8 show that while TabbyFlow performs slightly below TabSyn on MLE scores, it maintains competitive performance in all data sets. This presents an interesting contrast to earlier metrics, where TabbyFlow often outperformed TabSyn. The small performance gap suggests that MLE alone may not fully capture synthetic data quality, particularly given that TabbyFlow achieves comparable results with a substantially simpler architecture than TabSyn’s VAE-diffusion framework. On the other hand, we evaluate the performance of the models in preserving privacy as measured by DCR. The results, see in Table 7, show that Tabbyflow outperforms all other models on average, albeit marginally.
@@ Acknowledgments
This project was support by the Bosch Center for Artificial Intelligence. JWvdM additionally acknowledges support from the European Union Horizon Framework Programme (Grant agreement ID: 101120237). AGC acknowledges that the presentation of this paper at the conference was financially supported by the Amsterdam ELLIS Unit.
@@ A.2 Exponential Family VFM as Bregman Divergence
$\mathcal{L}_{\text{VFM}}(\theta)$ $=-\mathbb{E}_{t,x_{1},x}\left[\eta_{t}^{\theta}(x)\cdot\tau(x_{1})-A(\eta^{\theta}_{t}(x))\right]+\text{const.}$ (37)
$=-\mathbb{E}_{t,x_{1},x}\left[\mu^{\theta}_{t}(x)\cdot\eta^{\theta}_{t}(x)-A(\eta^{\theta}_{t}(x))+(\tau(x_{1})-\mu^{\theta}_{t}(x))\cdot\eta^{\theta}_{t}(x)\right]+\text{const.}$ (38)
$=-\mathbb{E}_{t,x_{1},x}\left[A^{*}(\mu_{t}^{\theta}(x))+(\tau(x_{1})-\mu_{t}^{\theta}(x))\cdot\eta_{t}^{\theta}(x)\right]+\text{const.}$ (39)
$=-\mathbb{E}_{t,x_{1},x}\left[-D_{A^{*}}\left(\tau(x_{1}),\mu_{t}^{\theta}(x)\right)+A^{*}(\tau(x_{1}))\right]+\text{const.}$ (40)
$=\mathbb{E}_{t,x_{1},x}\left[D_{A^{*}}\left(\tau(x_{1}),\mu_{t}^{\theta}(x)\right)\right]+\text{const.}$ (41)
$=\mathbb{E}_{t,x}\left[D_{A^{*}}\left(\mathbb{E}_{p_{t}(x_{1}\mid x)}[\tau(x_{1})],\mu_{t}^{\theta}(x)\right)\right]+\text{const.}$ (42)
$=\mathbb{E}_{t,x}\left[D_{A^{*}}\left(\mu_{t}(x),\mu_{t}^{\theta}(x)\right)\right]+\text{const.}$ (43)